# thanks: Corresponding authors:

### v1

@@ III.2.2 Timescale for stabilization
Beyond identifying the final steady state of the mutual repression system, it is important to characterize the time required for the system to reach steady state starting from different initial conditions. We define the relaxation time $\tau$ as the maximum of the times $\tau_{1}$ and $\tau_{2}$ taken by the two trajectories, $\bar{R}_{1}(\bar{t})$ and $\bar{R}_{2}(\bar{t})$, to reach $95\%$ of their respective steady states. Fig. 19(A) illustrates the influence of the initial condition on the final steady state in a symmetric system, where $\bar{K}=1$, $\omega_{1}=\omega_{2}=7.5$, and $c_{1}=c_{2}=10^{-6}\ \text{M}$. The figure reveals that the phase space is divided into two basins of attraction, each leading to one of the two stable steady states. The separatrix, plotted in green, denotes the boundary between these basins, and is derived analytically in Appendix J.
@@ IV Kinetics and Time Delays in Feed-Forward Loops
$\frac{dZ}{dt}=-\gamma Z+\frac{r_{0Z}+r_{1Z}\left(p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}+p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}\right)\textstyle+r_{2Z}\,\omega\,p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}{{1+p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}+p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}{+\omega\,p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}},$ (24)
for the regulation of Z by both X and Y. We assume here that the two proteins $Y$ and $Z$ have the same degradation rate $\gamma$. The production rates and dissociation constants are assumed to be different in general for each thermodynamic state. Specifically, $r_{iG}$ denotes the production rate of gene $G$ when $i$ number of transcription factor are bound. Further, $K_{G_{1}G_{2}}$ denotes the dissociation constant of transcription factor $G_{1}$ binding to gene $G_{2}$. $\omega$ is the cooperativity, which takes into account the extra interaction energy between $X$ and $Y$ when bound to the DNA. Finally, the probabilities $p_{\text{act}}^{X}(c_{X})$ and $p_{\text{act}}^{Y}(c_{Y})$ scale the activity of transcription factors $X$ and $Y$. Note also that we consider different effectors $c_{X}$ and $c_{Y}$ for the two genes that can be varied independently. These functions may be distinct in principle, and the analytic discussions here make no assumption regarding their nature. For numerical results, however, we assume these probability functions, and thus effector activity functions, to be the same regardless of target transcription factor for simplicity.
@@ IV.1 Characterizing delay responses in coherent feed-forward loops
Figure 21: Delay in coherent feed-forward loop response compared to simple regulation. (A) The input signal is applied by tuning $c_{X}$ as a step function: from high ($10^{-4}$ M) to low ($10^{-7}$ M) at $\bar{t}=6$, and back to high at $\bar{t}=18$. $c_{Y}$ is kept constant at low ($10^{-7}$ M) throughout. The effector concentration $c_{Y}$ is held constant at $10^{-7}$ M. These inputs determine the activation probabilities $p_{\text{act}}^{X}$ and $p_{\text{act}}^{Y}$, shown in red and purple, respectively. The second panel plots the time evolution of the dimensionless output concentration $\bar{Z}(\bar{t})$ under feed-forward and simple regulation schemes, with $\bar{r}_{0Y}=\bar{r}_{0Z}=0$, $\bar{r}_{1Y}=\bar{r}_{1Z}=2$, $\bar{r}_{2Z}=10$, $\bar{K}_{XZ}=\bar{K}_{YZ}=1$, and $\omega=1$. (B) Schematic demonstrating the two ways to equivalently quantify the delayed response of the feed-forward loop compared to simple regulation, captured by the shaded area between the two curves. One can either integrate over individual time delay measurements $\Delta\bar{t}(\bar{Z})$ as a function of $\bar{Z}$, or equivalently integrate the difference in responses $\Theta(\bar{t})$ as a function of $\bar{t}$.
@@
where $\bar{Z}_{i}$ is the initial steady state in the feed-forward setting before the change in $c_{X}$, and $\bar{Z}_{f}$ is the final steady state after the change that $\bar{Z}(\bar{t})$ relaxes to eventually. We derive Eqn. 30 explicitly from Eqns. 25 and  26 in Appendix K. Note that the sum of the first two terms in Eqn. 30 describes behavior of the same form as simple regulation in Eqn. 29. Therefore, by rescaling Eqn. 29 we can treat the exponential portion of Eqn. 30 as an equivalent $\bar{Z}_{\text{simple}}$ with the same relaxation dynamics as observed for simple regulation. We can also then express output response for the feed-forward loop as
@@
which encodes how the rates and cooperativity regulate output $\bar{Z}$ expression as a function of input signal $\mathcal{X}_{f}$. Interestingly, we note that all types of feed-forward loops have the same solution as Eqn. 32, except with a potentially different $\Phi$. Appendix K.4 discusses this in more detail. Finally, the quantity
@@
$\langle\Delta\bar{t}\rangle=\frac{\Phi(\bar{Z}_{f}-\bar{Z}_{i})^{-1}}{S(1+\omega\mathcal{X}_{f})}\log\Big(\frac{1+\mathcal{X}_{f}+\mathcal{Y}_{i}+\omega\mathcal{X}_{f}\mathcal{Y}_{i}}{S}\Big{missing}).$ (38)
@@
From Eqn. 38, we can analytically deduce whether a feed-forward loop delays or accelerates output response from the sign of $\langle\Delta\bar{t}\rangle$. In Appendix K.2, we prove that in general $\langle\Delta\bar{t}\rangle\leq 0$ for the coherent feed-forward loop, leading to delay for both the ON and OFF steps as seen in the example of Fig. 21(A).
@@ IV.2 Robustness of delayed response for different coherent logic gates
[…]
@@
We observe that the average delay $\langle\Delta\bar{t}\rangle$ in coherent feed-forward loops depends strongly on the dissociation constant $\bar{K}_{YZ}$, which sets the binding affinity of $Y$ to the promoter of $Z$—one of the interactions that distinguishes feed-forward loops from simple regulation. A clear trend emerges: ON steps (Fig. 22(A-C)(i)) exhibit stronger delays when $Y$ binding is weak (large $\bar{K}_{YZ}$), whereas OFF steps (Fig. 22(A-C)(ii)) show stronger delays when $Y$ binding is strong (small $\bar{K}_{YZ}$).
Since the feed-forward loop ultimately aims to control $Z$ activation, another relevant feature is the output amplitude $\Delta\bar{Z}=|\bar{Z}_{f}-\bar{Z}_{i}|$ for a given change in input $c_{X}$. A small $\Delta\bar{Z}$ would imply that $\bar{Z}$ remains nearly constant, making the response uninformative. However, we show in Appendix L that $\Delta\bar{Z}$ scales linearly with the production rates. Although the dependence of $\Delta\bar{Z}$ on other parameters is non-trivial, globally the effect of the rates dominates. We thus reserve a detailed discussion of the optimization of $\Delta\bar{Z}$ values for Appendix L and remain focused here on the average delay $\langle\Delta\bar{t}\rangle$.
@@ IV.3 Existence of pulse in incoherent feed-forward loop
is always negative. As a result, the average time difference $\langle\Delta\bar{t}\rangle$ as defined in Eqn. 37 is always positive, both for ON and OFF steps, meaning that the incoherent feed-forward loop accelerates the response for both transitions. In Appendix M, we remark that the definition and interpretation of $\langle\Delta\bar{t}\rangle$ is subtle when $\bar{Z}$ exhibits a pulse. Thus, we only consider $\langle\Delta\bar{t}\rangle$ when a pulse does not exist, and focus on the difference between the peak of the pulse and final steady state of $\bar{Z}$ for the pulses.
To further quantify this acceleration, we again compute the average time difference $\langle\Delta\bar{t}\rangle$ between the feed-forward loop and simple regulation trajectories with Eqn. 37. Fig. 24(A) shows $\langle\Delta\bar{t}\rangle$ across the phase space defined by $(\bar{K}_{XZ},\bar{K}_{YZ})$, when the output does not present a strong pulse. The definition of “strong” is described in Appendix M. We observe that acceleration is limited to a maximum value of 1. This upper bound arises because the most accelerated trajectory would consist of an instantaneous jump to the final steady state (green curves in Fig. 24(C) and (D)), corresponding to $\langle\Delta\bar{t}\rangle=1$ as shown in Appendix M. An example of a feed-forward trajectory that does not exhibit a strong pulse is shown in Fig. 24(C). The region of highest acceleration lies near the boundary separating the strongly pulsed and not strongly pulsed regimes. Interestingly, acceleration is only substantial in a restricted region of parameter space. For example, we see in Fig. 24(A) that high values of $\bar{K}_{YZ}$—corresponding to weak binding of $Y$—lead to negligible acceleration.
@@ IV.4 Continuous signal
Before proceeding to the analysis of the delay, we must first clarify the definition of simple regulation in the case of a continuous signal. We adopt the same definition as the step function case, where we fix concentration $\bar{Y}=1$ here. Further, in our previous analysis, when $c_{X}(\bar{t})$ is a step function, the choice of $\bar{Y}$ does not affect the dynamics of relaxation of the simple regulation output. However, when $c_{X}(\bar{t})$ is a continuous function, the choice of $\bar{Y}$ changes these dynamics slightly. Fortunately, the effect of $\bar{Y}$ is small and does not qualitatively change our observations in the rest of the section, allowing us to continue with the comparison between feed-forward loop and simple regulation (for a more detailed discussion, see Appendix N).
@@ V Discussion
Within this analysis of feed-forward loops, we demonstrate the crucial roles of effector concentration in our models. The leakiness of $p_{\text{act}}(c)$, for example, influences key metrics such as delay time. We also highlight how the sigmoidal shape of $p_{\text{act}}(c)$ enables a continuous change in effector concentration to be translated to a sharp signal when tuning the probability of transcription factors being active. Overall, while the dynamical behaviors of feed-forward loops can be rich, they are not always guaranteed. This emphasizes how behavior emerges from a delicate interplay of biochemical parameters rather than rigid circit logic alone, and underscores the need for further experimental and theoretical efforts toward understanding the functions and dynamics of feed-forward loops.
@@ Acknowledgements.
We have benefited enormously from conversations with Leonid Mirny, Ned Wingreen, Julie Theriot, Marc Kirschner, Jean Pierre Changeux, Jane Kondev, all of whom have helped us better understand the twin subjects brought together here, namely, genetic circuits and allosteric regulation of the macromolecules of the cell. We are grateful to the NIH for support through award numbers DP1OD000217 (Director’s Pioneer Award) and NIH MIRA 1R35 GM118043-01. This research is funded in part by the Gordon and Betty Moore Foundation GBMF12214 (doi.org/10.37807/GBMF12214) to RP and RJR as part of the grant on the “Listening to Molecules” project which has as its central mission to better understand the underlying mechanisms of allosteric regulation. RP is deeply grateful to the CZI Theory Institute Without Walls. HGG was supported by NIH R01 Awards R01GM139913 and R01GM152815, by the Koret-UC Berkeley-Tel Aviv University Initiative in Computational Biology and Bioinformatics, by a Winkler Scholar Faculty Award, and by the Chan Zuckerberg Initiative Grant CZIF2024-010479. H.G.G. is also a Chan Zuckerberg Biohub Investigator (Biohub – San Francisco).

### v2

@@ III.2.2 Timescale for stabilization
Beyond identifying the final steady state of the mutual repression system, it is important to characterize the time required for the system to reach steady state starting from different initial conditions. We define the relaxation time $\tau$ as the maximum of the times $\tau_{1}$ and $\tau_{2}$ taken by the two trajectories, $\bar{R}_{1}(\bar{t})$ and $\bar{R}_{2}(\bar{t})$, to reach $95\%$ of their respective steady states. Fig. 19(A) illustrates the influence of the initial condition on the final steady state in a symmetric system, where $\bar{K}=1$, $\omega_{1}=\omega_{2}=7.5$, and $c_{1}=c_{2}=10^{-6}\ \text{M}$. The figure reveals that the phase space is divided into two basins of attraction, each leading to one of the two stable steady states. The separatrix, plotted in green, denotes the boundary between these basins, and is derived analytically in Appendix K.
@@ IV Kinetics and Time Delays in Feed-Forward Loops
$\frac{dZ}{dt}=-\gamma Z+\frac{r_{0Z}+r_{1Z}\left(p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}+p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}\right)\textstyle+r_{2Z}\,\omega\,p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}{{1+p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}+p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}{+\omega\,p_{\text{act}}^{X}(c_{X})\frac{X}{K_{XZ}}p_{\text{act}}^{Y}(c_{Y})\frac{Y}{K_{YZ}}}}$ (24)
for the regulation of Z by both X and Y. We assume here that the proteins $Y$ and $Z$ have the same degradation rate $\gamma$. The production rates and dissociation constants are assumed to be different in general for each thermodynamic state. Specifically, $r_{iG}$ denotes the production rate of gene $G$ when $i$ number of transcription factor are bound. Further, $K_{G_{1}G_{2}}$ denotes the dissociation constant of transcription factor $G_{1}$ binding to gene $G_{2}$. $\omega$ is the cooperativity, which takes into account the extra interaction energy between $X$ and $Y$ when bound to the DNA. Finally, the probabilities $p_{\text{act}}^{X}(c_{X})$ and $p_{\text{act}}^{Y}(c_{Y})$ scale the activity of transcription factors $X$ and $Y$. Note also that we consider different effectors $c_{X}$ and $c_{Y}$ for the two genes that can be varied independently. These functions may be distinct in principle, and the analytic discussions here make no assumption regarding their nature. For numerical results, however, we assume these probability functions, and thus effector activity functions, to be the same regardless of target transcription factor for simplicity.
@@ IV.1 Characterizing delay responses in coherent feed-forward loops
Figure 21: Delay in coherent feed-forward loop response compared to simple regulation. (A) The input signal is applied by tuning $c_{X}$ as a step function: from high ($10^{-4}$ M) to low ($10^{-7}$ M) at $\bar{t}=6$, and back to high at $\bar{t}=18$. The effector concentration $c_{Y}$ is held constant at $10^{-7}$ M. These inputs determine the activation probabilities $p_{\text{act}}^{X}$ and $p_{\text{act}}^{Y}$, shown in red and purple, respectively. The second panel plots the time evolution of the dimensionless output concentration $\bar{Z}(\bar{t})$ under feed-forward and simple regulation schemes, with $\bar{r}_{0Y}=\bar{r}_{0Z}=0$, $\bar{r}_{1Y}=\bar{r}_{1Z}=2$, $\bar{r}_{2Z}=10$, $\omega=1$, and $\bar{K}_{XZ}=\bar{K}_{YZ}=1$. (B) Schematic demonstrating the two ways to equivalently quantify the delayed response of the feed-forward loop compared to simple regulation, captured by the shaded area between the two curves. One can either integrate over individual time delay measurements $\Delta\bar{t}(\bar{Z})$ as a function of $\bar{Z}$, or equivalently integrate the difference in responses $\Theta(\bar{t})$ as a function of $\bar{t}$.
@@
where $\bar{Z}_{i}$ is the initial steady state in the feed-forward setting before the change in $c_{X}$, and $\bar{Z}_{f}$ is the final steady state after the change that $\bar{Z}(\bar{t})$ relaxes to eventually. We derive Eqn. 30 explicitly from Eqns. 25 and 26 in Appendix L. Note that the sum of the first two terms in Eqn. 30 describes behavior of the same form as simple regulation in Eqn. 29. Therefore, by rescaling Eqn. 29 we can treat the exponential portion of Eqn. 30 as an equivalent $\bar{Z}_{\text{simple}}$ with the same relaxation dynamics as observed for simple regulation. We can also then express output response for the feed-forward loop as
@@
which encodes how the rates and cooperativity regulate output $\bar{Z}$ expression as a function of input signal $\mathcal{X}_{f}$. Interestingly, we note that all types of feed-forward loops have the same solution as Eqn. 32, except with a potentially different $\Phi$. Appendix L.4 discusses this in more detail. Finally, the quantity
@@
$\langle\Delta\bar{t}\rangle=\>\frac{1}{\bar{Z}_{f}-\bar{Z}_{i}}\int_{0}^{\infty}\Theta(\bar{t})d\bar{t}.$ (37)
@@
From Eqn. 38, we can analytically deduce whether a feed-forward loop delays or accelerates output response from the sign of $\langle\Delta\bar{t}\rangle$. In Appendix L.2, we prove that in general $\langle\Delta\bar{t}\rangle\leq 0$ for the coherent feed-forward loop, leading to delay for both the ON and OFF steps as seen in the example of Fig. 21(A).
@@ IV.2 Robustness of delayed response for different coherent logic gates
In Fig. 22, we show a parameter sweep for three sets of rates and cooperativity parameters that correspond to three different logic gates. For each set of parameters, both ON and OFF steps are studied. These calculations inspire several observations. First, for a given step undergone by $p_{\text{act}}^{X}(c_{X})$, we can computationally identify a maximum $\langle\Delta\bar{t}\rangle$ with respect to all other parameters. Further, this maximum is different for ON and OFF steps. For example, in Fig. 22(i), specific to the input step chosen here, we observe that in all logic gates the maximal $\langle\Delta\bar{t}\rangle$ achievable is about 1 for the ON step and 4.5 for the OFF step. To tie this back to units of time, we consider E. coli. Here, proteins tend to be stable over the timescale of a cell cycle and hence the dilution resulting from cell division becomes the effective degradation rate. Taking a rate of degradation of order $\gamma=10^{-2}\text{min}^{-1}$, we then get a maximum delay of $500\text{ min}$, which is more than $8\text{ hours}$. The asymmetry between maximal $\langle\Delta\bar{t}\rangle$ the ON and OFF steps resonates with the analytic discussion in the previous section. We show computational evidence in Appendix M that sweeps across all other parameters indicate the existence of a global maximum $\langle\Delta\bar{t}\rangle$. Note that the exact value and existence of this maximum is a result specific to function $p_{\text{act}}(c_{X})$ and allosteric parameters we chose to describe effector activity in Fig. 5.
@@
We observe that the average delay $\langle\Delta\bar{t}\rangle$ in coherent feed-forward loops depends strongly on the dissociation constant $\bar{K}_{YZ}$, which sets the binding affinity of $Y$ to the promoter of ZZ—one of the interactions that distinguishes feed-forward loops from simple regulation. A clear trend emerges: ON steps (Fig. 22(A-C)(i)) exhibit stronger delays when $Y$ binding is weak (large $\bar{K}_{YZ}$), whereas OFF steps (Fig. 22(A-C)(ii)) show stronger delays when $Y$ binding is strong (small $\bar{K}_{YZ}$).
Since the feed-forward loop ultimately aims to control $Z$ activation, another relevant feature is the output amplitude $\Delta\bar{Z}=|\bar{Z}_{f}-\bar{Z}_{i}|$ for a given change in input $c_{X}$. A small $\Delta\bar{Z}$ would imply that $\bar{Z}$ remains nearly constant, making the response uninformative. However, we show in Appendix M that $\Delta\bar{Z}$ scales linearly with the production rates. Although the dependence of $\Delta\bar{Z}$ on other parameters is non-trivial, globally the effect of the rates dominates. We thus reserve a detailed discussion of the optimization of $\Delta\bar{Z}$ values for Appendix M and remain focused here on the average delay $\langle\Delta\bar{t}\rangle$.
@@ IV.3 Existence of pulse in incoherent feed-forward loop
is always negative. As a result, the average time difference $\langle\Delta\bar{t}\rangle$ as defined in Eqn. 37 is always positive, both for ON and OFF steps, meaning that the incoherent feed-forward loop accelerates the response for both transitions. In Appendix N, we remark that the definition and interpretation of $\langle\Delta\bar{t}\rangle$ is subtle when $\bar{Z}$ exhibits a pulse. Thus, we only consider $\langle\Delta\bar{t}\rangle$ when a pulse does not exist, and focus on the difference between the peak of the pulse and final steady state of $\bar{Z}$ for the pulses.
To further quantify this acceleration, we again compute the average time difference $\langle\Delta\bar{t}\rangle$ between the feed-forward loop and simple regulation trajectories with Eqn. 37. Fig. 24(A) shows $\langle\Delta\bar{t}\rangle$ across the phase space defined by $(\bar{K}_{XZ},\bar{K}_{YZ})$, when the output does not present a strong pulse. The definition of “strong” is described in Appendix N. We observe that acceleration is limited to a maximum value of 1. This upper bound arises because the most accelerated trajectory would consist of an instantaneous jump to the final steady state (green curves in Fig. 24(C) and (D)), corresponding to $\langle\Delta\bar{t}\rangle=1$ as shown in Appendix N. An example of a feed-forward trajectory that does not exhibit a strong pulse is shown in Fig. 24(C). The region of highest acceleration lies near the boundary separating the strongly pulsed and not strongly pulsed regimes. Interestingly, acceleration is only substantial in a restricted region of parameter space. For example, we see in Fig. 24(A) that high values of $\bar{K}_{YZ}$—corresponding to weak binding of YY—lead to negligible acceleration.
@@ IV.4 Continuous signal
Before proceeding to the analysis of the delay, we must first clarify the definition of simple regulation in the case of a continuous signal. We adopt the same definition as the step function case, where we fix concentration $\bar{Y}=1$ here. Further, in our previous analysis, when $c_{X}(\bar{t})$ is a step function, the choice of $\bar{Y}$ does not affect the dynamics of relaxation of the simple regulation output. However, when $c_{X}(\bar{t})$ is a continuous function, the choice of $\bar{Y}$ changes these dynamics slightly. Fortunately, the effect of $\bar{Y}$ is small and does not qualitatively change our observations in the rest of the section, allowing us to continue with the comparison between feed-forward loop and simple regulation (for a more detailed discussion, see Appendix O).
@@ V Discussion
Within this analysis of feed-forward loops, we demonstrate the crucial roles of effector concentration in our models. The leakiness of $p_{\text{act}}(c)$, for example, influences key metrics such as delay time. We also highlight how the sigmoidal shape of $p_{\text{act}}(c)$ enables a continuous change in effector concentration to be translated to a sharp signal when tuning the probability of transcription factors being active. Overall, while the dynamical behaviors of feed-forward loops can be rich, they are not always guaranteed. This emphasizes how behavior emerges from a delicate interplay of biochemical parameters rather than rigid circuit logic alone, and underscores the need for further experimental and theoretical efforts toward understanding the functions and dynamics of feed-forward loops.
@@ Acknowledgements.
We have benefited enormously from conversations with Leonid Mirny, Ned Wingreen, Julie Theriot, Marc Kirschner, Jean Pierre Changeux, Jane Kondev, all of whom have helped us better understand the twin subjects brought together here, namely, genetic circuits and allosteric regulation of the macromolecules of the cell. We are grateful to the NIH for support through award numbers DP1OD000217 (Director’s Pioneer Award) and NIH MIRA 1R35 GM118043-01. This research is funded in part by the Gordon and Betty Moore Foundation GBMF12214 (doi.org/10.37807/GBMF12214) to RP and RJR as part of the grant on the “Listening to Molecules” project which has as its central mission to better understand the underlying mechanisms of allosteric regulation. RP is deeply grateful to the CZI Theory Institute Without Walls both for conversations with its members and support for the work itself. HGG was supported by NIH R01 Awards R01GM139913 and R01GM152815, by the Koret-UC Berkeley-Tel Aviv University Initiative in Computational Biology and Bioinformatics, by a Winkler Scholar Faculty Award, by the Chan Zuckerberg Initiative Grant CZIF2024-010479, and by the Miller Institute for Basic Research in Science, University of California Berkeley. H.G.G. is also a Chan Zuckerberg Biohub Investigator (Biohub – San Francisco).